OpenStudy (anonymous):

find the slope of a line (3,0) and (9,1)

OpenStudy (anonymous):

use the equation y2-y1/x2-x1

OpenStudy (anonymous):

m=y2-y1/x2-x1 m=(1-0)/(9-3) m=1/6 is your answer.

OpenStudy (bahrom7893):

rise over run, come on!

OpenStudy (anonymous):

sorry i put the wrong question. it was find the slope of the line (3,0) and (9,0)

OpenStudy (nenadmatematika):

it's zero

OpenStudy (anonymous):

so its undefined right?

OpenStudy (anonymous):

When ever your Y's are 0, the answer is 0

OpenStudy (anonymous):

Yes, to therfore its undefined

OpenStudy (anonymous):

it still says im wrong. but thank you

OpenStudy (anonymous):

The answer is 0/6 which is 0 IDK why it says you're wrong... (3,0) and (9,0) would mean both y2 and y1 are 0 so basically the slope will be 0 as well. Or you could just try inputting 0-0/9-3 = 0/6 lol

OpenStudy (anonymous):

Slope's 0 no problem. If it says wrong then try 0/6 if still no maybe the question is wrong or something's at fault in the system.